Flawless Cosmetic Company manufactures and distributes several different products. The company currently uses a plantwide allocation method for allocating overhead at a rate of $7 per direct labor hour. Loren is the department manager of the Makeup Department that produces Products - Concealer (C) and Glow Cream (GC). Jennifer is the department manager of the Hair Care Department that manufactures Product - Shampoo (S). The product costs (per case of 24 bottles) and other information are as follows:

Products
C GC S
Direct materials $ 100.00 $ 72.00 $ 48.00
Direct labor 42.00 31.50 12.00
Overhead 28.00 21.00 14.00
$ 170.00 $ 124.50 $ 74.00
Machine hours 4 2 3
Number of cases (per year) 300 500 600

Based on this information, if Flawless changes its allocation basis to machine hours, what is the overhead rate per machine hour?
Answer #1

total overhead = 28 + 21 + 14 = 63

total machine hour is 4+2+3 = 9

overhead rate per machine hour = total overhead cost / total machine hour

= 63 / 9

= 7 per machine hour